# Incentivize Contribution and Learn Parameters Too: Federated Learning with Strategic Data Owners

### v2

@@ 14             Send $d_{i}^{t+1}$ to the center
$\kappa\geqslant\max_{i}\left\{\frac{\displaystyle s^{\max}_{i}-s^{0}_{i}}{\displaystyle c\,\Delta_{i}}\right\},\qquad T_{0}>\left(\ln{\frac{f(w^{0},s^{\max})-f(w^{\texttt{{OPT}}},s^{\max})}{\epsilon}}\right)\Big{/}\left(\ln\frac{1}{{1-\frac{\nu}{M}}}\right)$ (9)
where, $f(w,s):=-\frac{1}{n}\sum_{i\in N}a_{i}(w,s)$ and $\Delta_{i}=\beta-\tau-\zeta$.
Remark (Strong convexity and smoothness). The convergence rate is $\mathcal{O}(\ln{\nicefrac{{1}}{{\epsilon}}})$ which is the same as that of Algorithm 1. However, Algorithm 2 is cleaner in terms of the phases of convergence, the choices of the step sizes $\gamma,\eta$. The structural assumptions like strong convexity and smoothness may be relaxed at the expense of weaker rates. For example, we observe convergence rates of $\mathcal{O}(\frac{1}{\sqrt{\epsilon}})$ for non-convex smooth functions and $\mathcal{O}(\frac{1}{\epsilon})$ for convex smooth functions.
Remark ($\gamma$ and rationality). Phase 1 of Algorithm 2 achieves the optimal contribution $s^{\max}$ without updating the learning parameter $w$. The speed of convergence of this phase depends on the parameter $\gamma=c$, which can be interpreted as the minimum rationality level of the society. For instance, if all agents are sufficiently rational, i.e., $c$ is large, they may converge to $s^{\max}$ in a single iteration. However, the algorithm converges even if agents are boundedly rational. In Phase 2 of the algorithm, the agents stop updating their contributions and only updates the gradients $d_{i}$ and the center accumulates them to update $w^{t}$.
Thanks to the smoothness and strong convexity properties of $f(\cdot,s^{\max})$, we can show that running Algorithm 2 guarantees that the model iterate $w^{T}$ converges to $w^{\texttt{{OPT}}}$ at an exponential speed as well.
Corollary 1 (Iterate Convergence).
With the same setup as above and $\gamma=c$ (universal constant), $\eta=\frac{2}{M+\nu}$, we obtain $\|w^{T}-w^{\texttt{{OPT}}}\|_{2}<\epsilon,\text{ and }s^{T}=s^{\max}$, where $T=\kappa+\tilde{T}_{0}$, with the same $\kappa$ as in Equation 9 and $\tilde{T}_{0}>\left(\ln{\frac{\|w^{0}-w^{\texttt{{OPT}}}\|_{2}}{\epsilon}}\right)/\left(\ln\frac{{1+\frac{\nu}{M}}}{{1-\frac{\nu}{M}}}\right)$.
5 Experiments
Our results in Sections 3 and 4 guarantee convergence of Algorithms 1 and 2 to a pure Nash equilibrium under the given assumptions. However, it is important to ask how these FL algorithms behave in practice with real datasets under more general setups. We capture results in these realistic settings in our experiments. We train models on the CIFAR-10  [19], FeMNIST, and Twitter datasets [3] and compare the social welfare of UPBReD, 2P-UPBReD, and FedAvg. We also consider another algorithm FedAvgStrategic. In this algorithm, the agents choose to contribute their data strategically as in UPBReD, and stick to this contribution in the entire training phase of FedAvg (details in Section C.1). The algorithm proposed in [25], which is closest to this paper, do not consider the simultaneous learning of parameters and strategy updates, and is therefore not meaningful to compare against.
The cost function used is linear $c_{i}(s_{i})=c_{i}s_{i}$, where $c_{i}$ is sampled from $U[0,1]$ for each agent. We choose our accuracy function to be $a_{i}(w^{k},s^{k})=r_{i}-\mathcal{L}_{i}(w^{k},s^{k})$ where the $\mathcal{L}_{i}(w^{k},s^{k})$ is the loss evaluated on agents $i$’s test dataset at the $k$-th round of training. The social welfare is given by $\sum_{i\in N}a_{i}(w^{k},s^{k})$. We perform strategy updates using the numerical derivative as follows:
The model parameters obtained by agent $i$ after performing one training step locally on the model initialized with $w^{k-1}$ is denoted by $w_{i}^{k}$. We use cross-entropy loss function for $\mathcal{L}$. All the experiments are performed on a NVIDIA RTX A5000 with a 24GB core and a NVIDIA RTX A6000 with a 48GB core. The experiments required to obtain the results we report take several days to run. For comparing social welfare against $\beta$ and number of agents, we randomize over 10 runs, sampling an integer $s^{0}_{i}$ from $U[\nicefrac{{s^{max}_{i}}}{{3}},\nicefrac{{2s^{max}_{i}}}{{3}}]$ and cost from $U[0,1]$. We plot the average value of time taken across $10$ runs.
For FeMNIST, we use $20\%$ of the dataset that generates users with data that is not identically distributed as per the codebase provided by [3]. We train a CNN with $6,603,710$ parameters for $100$ rounds. We choose $\gamma=0.1$ and $\beta=2$. We use the Adam optimizer with learning rate $0.001$.
For the Twitter dataset, we use $15\%$ of the dataset that generates users with data that is identically distributed as per the codebase provided by [3]. We further group $1000$ agents together to create agents with larger data shares. We use an LSTM network to train for $50$ rounds. We choose $\gamma=0.5$ and $\beta=2$. We use the Adam optimizer with learning rate $0.00001$.
All results are summarized in Figure 1. We provide more details of the model architecture for all models in Section C.2.
(a) CIFAR-10
(b) FeMNIST
(c) Twitter
We observe that 2P-UPBReD and FedAvg have the same social welfare performance w.r.t. the number of agents (first column) and running times are nearly the same for all the datasets (third column). 2P-UPBReD outperforms UPBReD in nearly all the runs with respect to social welfare. We also observe that 2P-UPBReD converges to the maximum contribution of the agents for values of $\beta$ that are lower than the theoretically predicted ones. We also note that the data contributors get paid and data consumers pay in 2P-UPBReD and their payments are monotone non-decreasing with data contribution (Figure 2). The utilities vary because the dataset on which the model is tested to obtain the accuracy varies for the agents.
6 Conclusions and future work
In this paper, we proposed UPBReD, which ensures convergence to a Nash equilibrium along with the model parameters $w$. However, $w$ may deviate from the optimal $w^{\texttt{{OPT}}}$, so we introduced a two-phase mechanism, 2P-UPBReD, that guarantees full agent contribution and convergence to $w^{\texttt{{OPT}}}$. This comes at the cost of requiring monetary transfers, though these can be internal among agents.
An immediate direction for future work is to incorporate budget constraints for agents, limiting the maximum payment each can make while preserving approximate guarantees. Another important extension is to handle heterogeneous data distributions, allowing agents to form clusters and learn optimal parameters within each cluster.
Acknowledgments
This work has been primarily supported by an Amazon grant (RD/0123-AMAZO09-005) and AG likes to acknowledge that. Moreover, SN likes to also acknowledge the supports of a Core Research Grant (CRG/2023/001442) from SERB, Govt. of India, a TCS Research grant (RD/0123-TCSIL18-001), and an IIT Bombay grant (2022058).
Appendix A Appendix to Section 3
A.1 Proof of Theorem 1
$\quad+{H(w^{\prime},s^{t},\mu^{t})(w^{t+1}-w^{t})},$
$\quad+\tilde{H}(w^{t},s^{\prime})(s^{t+1}-s^{t}),$
Using the update rules of Algorithm 1 given by Equation 3 and Equation 4, we get
$\quad+{H(w^{\prime},s^{t},\mu^{t})\eta\tilde{g}(w^{t},s^{t})}$
$\quad+\tilde{H}(w^{t},s^{\prime})\gamma g(w^{t},s^{t},\mu^{t})$
Using triangle inequality on each of these identities, we get
[…]
From condition 1 of 1, we get that $v^{\top}(G+\lambda I_{n\times n})v\leqslant 0,\forall v\in\mathbb{R}^{n}$ and $v^{\prime\top}(\tilde{G}+\lambda I_{m\times m})v^{\prime}\leqslant 0,\forall v^{\prime}\in\mathbb{R}^{m}$. In particular, for $v=g(w^{t},s^{t},\mu^{t})$ and $v^{\prime}=\tilde{g}(w^{t},s^{t})$, we get
$\begin{split}g(w^{t},s^{t},\mu^{t})^{\top}G(w^{t},s^{\prime},\mu^{t})g(w^{t},s^{t},\mu^{t})&\leqslant-\lambda\|g(w^{t},s^{t},\mu^{t})\|_{2}^{2}\\ \tilde{g}(w^{t},s^{t})^{\top}\tilde{G}(w^{\prime},s^{t})\tilde{g}(w^{t},s^{t})&\leqslant-\tilde{\lambda}\|\tilde{g}(w^{t},s^{t})\|_{2}^{2}\end{split}$ (12)
Consider the square of the first term of the RHS of the inequality of $g$ in Equation 11
$\|(I_{n\times n}+\gamma.G(w^{t},s^{\prime},\mu^{t}))g(w^{t},s^{t},\mu^{t})\|^{2}$\mbox{}\hfil\\ &=\|g(w^{t},s^{t},\mu^{t})\|_{2}^{2}+\gamma^{2}.\|G(w^{t},s^{t})g(w^{t},s^{t},\mu^{t})\|_{2}^{2}\\ &\quad+2\gamma g(w^{t},s^{t},\mu^{t})^{\top}G(w^{t},s^{\prime},\mu^{t})g(w^{t},s^{t},\mu^{t})\\ &\leqslant\|g(w^{t},s^{t},\mu^{t})\|_{2}^{2}+\gamma^{2}.n^{2}L^{2}\|g(w^{t},s^{t},\mu^{t})\|_{2}^{2}-2\gamma\lambda\|g(w^{t},s^{t},\mu^{t})\|_{2}^{2}\\ &=(1+\gamma^{2}.n^{2}L^{2}-2\gamma\lambda)\|g(w^{t},s^{t},\mu^{t})\|_{2}^{2}\end{split}$ (13)
where the equality comes by expanding the squared norm and the inequality comes from the facts that (i) $\|G(w^{t},s^{t})g(w^{t},s^{t},\mu^{t})\|_{2}^{2}\leqslant\|G(w^{t},s^{t})\|_{F}^{2}\|g(w^{t},s^{t},\mu^{t})\|_{2}^{2}$, where $\|A\|_{F}:=\sqrt{\sum_{i}\sum_{j}|A_{ij}|^{2}}$ is the Frobenius norm of a matrix $A$ and (ii) using Equation 12. By condition 2 of 1, $\|G(w^{t},s^{t})\|_{F}^{2}\leqslant n^{2}L^{2}$. Hence, we get
$\|(I_{n\times n}+\gamma.G(w^{t},s^{t}))g(w^{t},s^{t},\mu^{t})\|$
$\leqslant\sqrt{1+\gamma^{2}.n^{2}L^{2}-2\gamma\lambda}\cdot\|g(w^{t},s^{t},\mu^{t})\|_{2}$
given the term inside the square root is positive.
Consider the second term of the RHS of the inequality of $g$ in Equation 11, where
$\|H(w^{\prime},s^{t},\mu^{t})\tilde{g}(w^{t},s^{t})\|_{2}^{2}$ $\leqslant\|H(w^{\prime},s^{t},\mu^{t})\|_{op}^{2}\|\tilde{g}(w^{t},s^{t})\|_{2}^{2}$
$\leqslant P^{2}\|\tilde{g}(w^{t},s^{t})\|_{2}^{2}$
by condition 2 of 1.
Defining $\alpha:=1+\gamma^{2}.n^{2}L^{2}-2\gamma\lambda,\beta:=P^{2}\eta^{2},\tilde{\alpha}:=1+\eta^{2}.m^{2}\tilde{L}^{2}-2\eta\tilde{\lambda},\tilde{\beta}:=\tilde{P}^{2}\gamma^{2}$, and carrying out a similar analysis for $\tilde{g}$ in Equation 11, we get
$\begin{split}\|g(w^{t+1},s^{t+1},\mu^{t+1})\|_{2}&\leqslant\sqrt{\alpha}\|g(w^{t},s^{t},\mu^{t})\|_{2}+\sqrt{\beta}\|\tilde{g}(w^{t},s^{t})\|_{2}\\ \|\tilde{g}(w^{t+1},s^{t+1})\|_{2}&\leqslant\sqrt{\tilde{\alpha}}\|\tilde{g}(w^{t},s^{t})\|_{2}+\sqrt{\tilde{\beta}}\|g(w^{t},s^{t},\mu^{t})\|_{2}\end{split}$ (14)
Adding the inequalities of Equation 14
$\|g(w^{t+1},s^{t+1},\mu^{t+1})\|_{2}+\|\tilde{g}(w^{t+1},s^{t+1})\|_{2}$\mbox{}\hfil\\ &\leqslant\max\{\sqrt{\alpha}+\sqrt{\tilde{\beta}},\sqrt{\tilde{\alpha}}+\sqrt{\beta}\}\left(\|g(w^{t},s^{t},\mu^{t})\|_{2}+\|\tilde{g}(w^{t},s^{t})\|_{2}\right)\end{split}$ (15)
To ensure that the above inequality is a contraction, we need to ensure $\sqrt{\alpha}+\sqrt{\tilde{\beta}},\sqrt{\tilde{\alpha}}+\sqrt{\beta}\in(0,1)$. These imply (i) $10<\alpha<1,0<\tilde{\alpha}<$ 1, (ii) $\sqrt{\tilde{\beta}}<1,\sqrt{\beta}<1$, and (iii) $\sqrt{\alpha}+\sqrt{\tilde{\beta}}<1,\sqrt{\tilde{\alpha}}+\sqrt{\beta}<1$. We can solve for $\gamma$ and $\eta$ from these inequalities and obtain the sufficient conditions when $\lambda>\tilde{P}$ and $\tilde{\lambda}>P$:
Notice that from conditions 1 and 2 of 1, we have that $\lambda<L$ and $\tilde{\lambda}<\tilde{L}$. Thus $n^{2}L^{2}-\tilde{P}^{2}>0$ and $m^{2}\tilde{L}^{2}-P^{2}>0$. For the $\gamma,\eta$ chosen as above, we find $W_{1}=\sqrt{\alpha}+\sqrt{\tilde{\beta}}<1$ and $W_{2}=\sqrt{\tilde{\alpha}}+\sqrt{\beta}<1$, and hence $W=\max\{W_{1},W_{2}\}<1$. Therefore,
$\|g(w^{t+1},s^{t+1},\mu^{t+1})\|_{2}+\|\tilde{g}(w^{t+1},s^{t+1})\|_{2}$
Recursively iterating over this inequality, we get
$\|g(w^{T},s^{T},\mu^{T})\|_{2}+\|\tilde{g}(w^{T},s^{T})\|_{2}$
$\leqslant W^{T}\left(\|{g}(w^{0},s^{0},\mu^{0})\|_{2}+\|\tilde{g}(w^{0},s^{0})\|_{2}\right).$
Defining $E=\|g(w^{0},s^{0},\mu^{0})\|_{2}+\|\tilde{g}(w^{0},s^{0})\|_{2}$ and
$T_{0}(w^{0},s^{0})=\nicefrac{{\left(\ln{\frac{\displaystyle E}{\displaystyle\epsilon}}\right)}}{{\left(\ln{\frac{\displaystyle 1}{\displaystyle W}}\right)}},$
we get that for all $T\geqslant T_{0}(w^{0},s^{0})$,
$\|g(w^{T},s^{T},\mu^{T})\|_{2}+\|\tilde{g}(w^{T},s^{T})\|_{2}<\epsilon$
This completes the proof.
A.2 Feasible region of $\gamma$ for Theorem 1 to hold
Figure 3 shows feasible regions of $\gamma$ for certain $n,L,\lambda,\tilde{P}$.
(a) $\tilde{P}=1$
(b) $\lambda=1.1$
A.3 Application of Theorem 1: effective costs are only personal
One special but practical case of the above setup is when the effective cost is only borne by the agent. Mathematically, it is represented as $z_{i}(s_{i},s_{-i})=c_{i}(s_{i}),\forall s_{i}\in S_{i},s_{-i}\in S_{-i},\forall i\in N$. We assume convex cost functions $c_{i},\forall i\in N$, which is a commonly used assumption [20] in the literature. This setup is consistent with the assumptions of 1, and according to Theorem 1, Algorithm 1 converges to a Nash equilibrium, which we denote as $(w^{*},s^{*})$. However, it is possible that none of the following things happen: (i) $s^{*}\neq s^{\max}$, i.e., the agents do not contribute their entire data for the federated learning process, leading to a suboptimal learning, or (ii) $w^{*}\neq w^{\texttt{{OPT}}}$ (see Equation 2 for the definition), which is neither the objective of the center nor the agents. The following example shows such an instance.
Example 1 (NE different from socially optimal).
Consider the federated learning setup with two agents that have linear costs and are trying to learn a model with two parameters. The accuracy function is identical and is given by $a_{i}(w,s)=1-\frac{(1-w_{1})^{2}+(2-w_{2})^{2}}{s_{1}+s_{2}},i=1,2$. The cost functions for agents $1$ and $2$ are given by $c_{1}(s_{1})=0.04\cdot s_{1}$ and $c_{2}(s_{2})=0.02\cdot s_{2}$ respectively. The agents’ strategy sets are $S_{i}=[0,5],i=1,2$. Notice that, at $s=s^{\max}=(5,5)$, the value of $w^{\texttt{{OPT}}}$ (see Equation 2) is $(1,2)$. It yields an optimal social welfare of $2$. However, this is not an NE, since the derivatives $\partial u_{i}/\partial s_{i}|_{s_{i}^{\max}}<0$ for both $i=1,2$. Hence, the best response of each agent is to reduce $s_{i}$. However, from this point, with the choice of $\gamma=0.25$ and $\eta=0.25$, Algorithm 1 converges to an NE profile of $w^{*}=(0.5,1.5),s^{*}=(0,5)$ that yields a social welfare of $1.8$.
Appendix B Appendix to Section 4
B.1 Proof of Lemma 1
@@
From Equation 8, we get the derivative of the utility function for $s_{i}\in(0,s_{i}^{\max})$ (note that this holds only for the interior of $S_{i}$, the derivative at the boundaries are zero by definition of function $g$ in Section 3) to be $[g(w,s,\mu)]_{i}=\frac{\partial}{\partial s_{i}}u_{i}(w,s)=\frac{\partial}{\partial s_{i}}a_{i}(w,s)-c_{i}^{\prime}(s_{i})+\beta$, where $c_{i}^{\prime}(s_{i})=\frac{d}{ds_{i}}c_{i}(s_{i})$. Using Assumption 2 and choosing in Equation 7 the parameter $\beta>\zeta+\tau$, we have
$[g(w,s,\mu)]_{i}$ $=\frac{\partial}{\partial s_{i}}a_{i}(w,s)-c_{i}^{\prime}(s_{i})+\beta\geqslant-\tau-\zeta+\beta>0.$
The last inequality holds since $\beta>\zeta+\tau$, $\forall i\in N$. ∎
B.2 Proof of Theorem 2
@@
$\begin{split}[g(w^{0},s^{t},\mu^{t})]_{i}\geqslant-\tau-\zeta+\beta>0\end{split}$ (16)
For the choice of $\beta>\zeta+\tau$ , we get $\Delta_{i}:=\beta-\zeta-\tau>0$. Applying the update rule for $s$ given by Algorithm 2 in phase 1, we get
$s^{t+1}$ $=s^{t}+\gamma g(w^{0},s^{t},\mu^{t})\geqslant s^{t}+\gamma\Delta_{i},$
@@
$s^{t}_{i}\geqslant s^{0}_{i}+\gamma(t\Delta_{i}).$
@@
$s^{t}_{i}$ $\geqslant s^{\max}_{i}-l_{i}+\gamma(t\Delta_{i})$
Substituting $t=\kappa$ where $\kappa\geqslant\frac{\displaystyle l_{i}}{\displaystyle\Delta_{i}\gamma}$, we obtain
$\begin{split}s^{\kappa}_{i}&\geqslant s^{\max}_{i}-l_{i}+\gamma\Delta_{i}(\frac{\displaystyle l_{i}}{\displaystyle\Delta_{i}\gamma})\\ &\geqslant s^{\max}_{i}-l_{i}+l_{i}\\ &\geqslant s^{\max}_{i}\\ s_{i}^{\kappa}&\geqslant s_{i}^{\max}\end{split}$ (17)
@@
Phase 2
For $t>\kappa$, we can focus on the second phase of the algorithm. From the fixed point property of $s^{\max}$, we can now focus entirely on the function $f(w,s^{\max})=-\frac{1}{n}\sum_{i\in N}a_{i}(w,s^{\max})$. From the algorithmic description, the second phase is just a simple gradient descent, run by the center, on $f(w,s^{\max})$. This is easy to see since at every iteration the center gets $d_{i}^{t+1}$ from all the agents $i\in N$, aggregates them and construct $\tilde{g}(w^{t},s^{\max})$, which is the gradient of $f(w,s^{\max})$ computed at $w^{t}$. We denote $f(w,s^{\max})$ with $f(w)$ in this part for notational cleanliness.
Note that the gradient descent is run with initialization $w^{\kappa}$ (which is same as $w^{0}$ as given by first phase of the algorithm). We exploit the strong convexity and smoothness of $f(w,s^{\max})$. Running the second phase for $T_{0}$ iterations, using [35], we obtain
$f(w^{T_{0}})-f(w^{\texttt{{OPT}}})\leq(1-\frac{\nu}{M})^{T_{0}}[f(w^{0})-f(w^{\texttt{{OPT}}})].$
@@
which proves the theorem.
B.3 Proof of Corollary 1
@@
The proof of this follows in the same lines as Theorem 2. Using the choice of $\beta$, we first ensure that with $\kappa$ steps, we obtain $s_{i}^{\kappa}=s^{\max}_{i}$ for all $i\in N$. Now, the framework is same as minimization of a strongly convex and smooth function $f(w)$ with initialization $w^{\kappa}=w^{0}$. Using [35], with $\eta=\frac{2}{M+\nu}$, we obtain the iterate convergence, namely $\|w^{\tilde{T}_{0}}-w^{\texttt{{OPT}}}\|_{2}\leqslant\left(\frac{1-\frac{\nu}{M}}{1+\frac{\nu}{M}}\right)^{\tilde{T}_{0}}\|w^{0}-w^{\texttt{{OPT}}}\|_{2}$. Taking log both sides implies the result. ∎
Appendix C Appendix to Section 5
C.1 Algorithm for FedAvgStrategic
In the FedAvgStrategic algorithm, the agents choose to contribute their data strategically as in UPBReD, and stick to this contribution in the entire training phase of FedAvg. FedAvgStrategic is formally given by Algorithm 3.

### v3

@@ 14      Send $d_{i}^{t+1}$ to the center
$\kappa\geqslant\max_{i}\left\{\frac{\textstyle s^{\max}_{i}-s^{0}_{i}}{\textstyle c\,\Delta}\right\},$ (16)
$T_{0}>\left(\ln{\frac{f(w^{0},s^{\max})-f(w^{\texttt{{OPT}}},s^{\max})}{\epsilon}}\right)\Big/\left(\ln\frac{1}{{1-\frac{\nu}{M}}}\right)$ (17)
where, $f(w,s):=-\frac{1}{n}\sum_{i\in N}v_{i}(w,s)$ and $\Delta=\beta-\tau-\zeta$.
@@
$\begin{split}\textstyle[g(w^{0},s^{t},\mu^{t})]_{i}\geqslant-\tau-\zeta+\beta>0\end{split}$ (18)
For the choice of $\beta>\zeta+\tau$ , we get $\Delta:=\beta-\zeta-\tau>0$. Applying the update rule for $s$ given by Algorithm˜3 in phase 1, we get
$s^{t+1}_{i}$ $=s^{t}_{i}+\gamma[g(w^{0},s^{t},\mu^{t})]_{i}\geqslant s^{t}_{i}+\gamma\Delta,$
@@
$s^{t}_{i}\geqslant s^{0}_{i}+\gamma(t\Delta).$
@@
$s^{t}_{i}$ $\geqslant s^{\max}_{i}-l_{i}+\gamma(t\Delta)$
Substituting $t=\kappa$ where $\kappa\geqslant\frac{\textstyle l_{i}}{\textstyle\Delta\gamma}$, we obtain
$\begin{split}\textstyle s^{\kappa}_{i}&\textstyle\geqslant s^{\max}_{i}-l_{i}+\gamma\Delta(\frac{\textstyle l_{i}}{\textstyle\Delta\gamma})\\ \textstyle&\textstyle\geqslant s^{\max}_{i}-l_{i}+l_{i}\\ \textstyle&\textstyle\geqslant s^{\max}_{i}\\ \textstyle s_{i}^{\kappa}&\textstyle\geqslant s_{i}^{\max}\end{split}$ (19)
@@
Phase 2 For $t>\kappa$, we can focus on the second phase of the algorithm. From the fixed point property of $s^{\max}$, we can now focus entirely on the function $f(w,s^{\max})=-\frac{1}{n}\sum_{i\in N}v_{i}(w,s^{\max})$. From the algorithmic description, the second phase is just a simple gradient descent, run by the center, on $f(w,s^{\max})$. This is easy to see since at every iteration the center gets $d_{i}^{t+1}$ from all the agents $i\in N$, aggregates them and construct $\tilde{g}(w^{t},s^{\max})$, which is the gradient of $f(w,s^{\max})$ computed at $w^{t}$. We denote $f(w,s^{\max})$ with $f(w)$ in this part for notational cleanliness.
Note that the gradient descent is run with initialization $w^{\kappa}$ (which is same as $w^{0}$ as given by first phase of the algorithm). We exploit the strong convexity and smoothness of $f(w,s^{\max})$. Running the second phase for $T_{0}$ iterations, using Wright and Recht (2022), we obtain
$f(w^{T_{0}})-f(w^{\texttt{{OPT}}})\leqslant(1-\frac{\nu}{M})^{T_{0}}[f(w^{0})-f(w^{\texttt{{OPT}}})].$
@@
which proves the theorem. Remark (Strong convexity and smoothness). The convergence rate is $\mathcal{O}(\ln{\nicefrac{{1}}{{\epsilon}}})$ which is the same as that of Algorithm˜1. However, Algorithm˜3 is cleaner in terms of the phases of convergence, the choices of the step sizes $\gamma,\eta$. The structural assumptions like strong convexity and smoothness may be relaxed at the expense of weaker rates. For example, we observe convergence rates of $\mathcal{O}(\frac{1}{\sqrt{\epsilon}})$ for non-convex smooth functions and $\mathcal{O}(\frac{1}{\epsilon})$ for convex smooth functions. ∎
Remark ($\gamma$ and rationality). Phase 1 of Algorithm˜3 achieves the optimal contribution $s^{\max}$ without updating the learning parameter $w$. The speed of convergence of this phase depends on the parameter $\gamma=c$, which can be interpreted as the minimum rationality level of the society. For instance, if all agents are sufficiently rational, i.e., $c$ is large, they may converge to $s^{\max}$ in a single iteration. However, the algorithm converges even if agents are boundedly rational. In Phase 2 of the algorithm, the agents stop updating their contributions and only update the gradients $d_{i}$ and the center accumulates them to update $w^{t}$.
Thanks to the smoothness and strong convexity properties of $f(\cdot,s^{\max})$, we can show that running Algorithm˜3 guarantees that the model iterate $w^{T}$ converges to $w^{\texttt{{OPT}}}$ at an exponential speed as well.
Corollary 1 (Iterate Convergence).
With the same setup as above and $\gamma=c$ (universal constant), $\eta=\frac{2}{M+\nu}$, we obtain $\|w^{T}-w^{\texttt{{OPT}}}\|_{2}<\epsilon,\text{ and }s^{T}=s^{\max}$, where $T=\kappa+\tilde{T}_{0}$, with the same $\kappa$ as in Equation˜16 and $\tilde{T}_{0}>\left(\ln{\frac{\|w^{0}-w^{\texttt{{OPT}}}\|_{2}}{\epsilon}}\right)/\left(\ln\frac{{1+\frac{\nu}{M}}}{{1-\frac{\nu}{M}}}\right)$.
@@
The proof of this follows in the same lines as Theorem 3. Using the choice of $\beta$, we first ensure that with $\kappa$ steps, we obtain $s_{i}^{\kappa}=s^{\max}_{i}$ for all $i\in N$. Now, the framework is same as minimization of a strongly convex and smooth function $f(w)$ with initialization $w^{\kappa}=w^{0}$. Using Wright and Recht (2022), with $\eta=\frac{2}{M+\nu}$, we obtain the iterate convergence, namely $\|w^{\tilde{T}_{0}}-w^{\texttt{{OPT}}}\|_{2}\leqslant\left(\frac{1-\frac{\nu}{M}}{1+\frac{\nu}{M}}\right)^{\tilde{T}_{0}}\|w^{0}-w^{\texttt{{OPT}}}\|_{2}$. Taking log both sides implies the result. ∎
Remark (Truthfulness). An agent’s true type is $(s_{i}^{max},d_{i}^{t})$ in every iteration of phase $2$. We can use the payment scheme $p_{i}^{*}$ as defined in Equation˜13, with the necessary assumptions, in addition to the $z_{i}(s_{i},s_{-i})$ defined in this section. This modification will make the mechanism truthful with the trade-off that it will no longer be budget-balanced.
Remark (Robustness to adversarial behavior). In 2P-UPBReD, we have assumed that the agents report their gradients truthfully in phase $2$. If the agents misreport by sending arbitrary gradients not computed on any data, we have discussed one option of enforcing additional payment in the previous remark that compromises budget balance. If we do not wish to enforce additional payments, we can use some established methods in literature that are robust to adversarial agents (Yin et al., 2018; Karimireddy et al., 2020a; Ghosh et al., 2021). We describe one such method called trimmed mean  (Yin et al., 2018), in Section˜4.2 in the supplementary material and show how this method performs compared to UPBReD and 2P-UPBReD in the presence of adversarial agents in Section˜5.
4.2 Adversarial agents
Our results for 2P-UPBReD assume that agents are honest and report their gradients truthfully. However, the agents may be adversaries who misreport their gradients. Prior works (Yin et al., 2018) describe some simple ways to deal with this. When the ratio of adversarial agents in the system is low we use 2P-UPBReD along with a trimmed mean algorithm described in (Yin et al., 2018).
Let $\alpha$ be the fraction of adversarial agents in the system. Phase $1$ of the algorithm remains the same with agents first converging to $s^{\max}$. In the second phase which is a classic FedAvg algorithm the agents only report gradients computed on their training data. Instead of aggregating all gradients, for each coordinate of the gradients the values of the gradient corresponding to that coordinate are sorted. The lowest and highest $\alpha$ fraction of these values are discarded and the remaining gradient coordinates are used to update the parameter. We detail the algorithm in Algorithm˜4. In Section˜5 we show that this algorithm is robust to adversaries when the fraction of adversarial agents is small.
1
[…]
2
Output: $w^{T}$
Center broadcasts $w^{0},s^{0}$, set $t=0$
/* Begin phase 1 */
3 while $s\neq s^{\max}$ do
4 Center broadcasts $s^{t}$
5 for agent $i\in N$ in parallel do
6    $s_{i}^{t+1}=s_{i}^{t}+\gamma[g(w^{0},s^{t},\mu^{t})]_{i}$
7      Send $s_{i}^{t+1}$ to the center
8
 $t=t+1$
 /* End phase 1 */
9
/* Begin phase 2 */
10 for $t=0$ to $T$ do
11 for agent $i\in N$ in parallel do
12      Compute local gradient: $d_{i}^{t+1}=\nabla_{w}\>v_{i}(w^{t},s_{i}^{\max},s_{-i}^{\max})$
13      Send $d_{i}^{t+1}$ to the center
14
15
16 foreach $j\in[m]$ do
17      Construct $W_{j}=\{d_{ij}|i\in N\}$.
18      Sort $W_{j}$ into an ascending array $C$.
19    $C=C[n\alpha,n-n\alpha]$
20    $W_{j}=\{d_{ij}|d_{ij}\in C\}$
21    Center updates: $w^{t+1}_{j}=w^{t}_{j}+\frac{\eta}{n}\sum_{d_{ij}\in W_{j}}\tilde{d}_{ij}^{t+1}$
22
23 Center broadcasts $w^{t+1}=w$
24
return $w^{T}$
/* End phase 2 */
Algorithm 4 Trimmed mean for 2P-UPBReD (trim-2P-UPBReD)
5 Experiments
Our results in Sections˜3 and 4 guarantee convergence of Algorithms˜1 and 3 to a pure Nash equilibrium under the given assumptions. However, it is important to ask how these FL algorithms behave in practice with real datasets under more general setups. We capture results in these realistic settings in our experiments. We train models on the CIFAR-10  (Krizhevsky, 2009), FEMNIST, and Twitter datasets (Caldas et al., 2019) and compare the social welfare of UPBReD, 2P-UPBReD, and FedAvg. We also consider another algorithm FedAvgStrategic. In this algorithm, the agents choose to contribute their data strategically as in UPBReD, and stick to this contribution in the entire training phase of FedAvg (details in Algorithm˜5). The algorithm proposed in (Murhekar et al., 2023), which is closest to this paper, do not consider the simultaneous learning of parameters and strategy updates, and is therefore not meaningful to compare against.